A ball is launched horizontally from the top of a 12 m tall vertical cliff. The ball has an initial

horizontal speed of 8.4 m/s.


1
Expert's answer
2021-10-12T12:35:30-0400

A ball is launched horizontally from the top of a 12 m tall vertical cliff. The ball has an initial

horizontal speed of 8.4 m/s. What is the flight time of the ball?


"h=gt^2/2\to t=\sqrt{2h/g}=\sqrt{2\cdot12/9.8}=1.56\ (s)"